5g of Zinc displaces 4.846 g of Copper from CuSO4. If Zinc has Equivalent Weight of 32.5. Find the Equivalent Weight of Copper.
 

Grade:11

2 Answers

Vikas TU
4 years ago
Weight of copper liberated/weight of zn rected = equivalent weight of copper / eq weight of Zn 
so, 1/5g = Eq. weight of copper/ 32.5
Eq. weight of copper = 32.5/5 =6.5 
Hope it helps.
Pragnya
4 years ago
Equation made is:
Zn+CuSO4----->Cu+ZnSO4
Eq. Weight/Weight given of Zn=Eq. weight/Weight given of Cu
32.5/5=Eq weight/4.845
Eq Weight of Cu=6.5*4.845=31.5(aprrox~)